The following table shows the actual demand observed over the last 4 years: 

Year1234
Demand88410

Using exponential smoothing with a = 0.40 and a forecast for year 1 of 7.0, provide the forecast from periods 2 through 5 (round your responses to one decimal place)

Year12345
Forecast (ES)7.0



Provide the forecast from periods 2 through 5 using the naive approach (enter your responses as whole numbers) 

Year2345
Forecast (NA)
Answer #1

Using the exponential smoothing method

F(2) = 0.4*Actual(1) + (1-0.4)*F(1)

= 0.4*8 + 0.6*7

= 7.4

F(3) = 0.4*8 + 0.6*7.4

= 7.6

F(4) = 0.4*4 + 0.6*7.6

= 6.2

F(5) = 0.4*10 + 0.6*6.2

= 7.7

Using the naive forecasting method

F(2) = Actual(1)

= 8

F(3) = Actual(2)

= 8

F(4) = 4

F(5) = 10

Please provide up votes, if you are satisfied with the response. Comment in case of any issue regarding the same